OpenStudy (anonymous):

If m∠4 = (3x)° and m∠8 = (x + 40)°, what is the measure of ∠4?

OpenStudy (nali):

is ther is a picture with this question

OpenStudy (anonymous):

No, sorry

OpenStudy (nali):

ok are these angles equal?

OpenStudy (anonymous):

Uhm...I don't know

OpenStudy (nali):

is that a hw question i mean is there is a source with u like the lesson before that question that can help give more info to this question

OpenStudy (anonymous):

Yes it's a homework question. The lesson was "angle relationships and parallel lines". I can't link, sorry. :(

OpenStudy (nali):

ok

OpenStudy (anonymous):

draw it

OpenStudy (anonymous):

The question? I'm not sure how to

OpenStudy (anonymous):

see the 3 blue buttons below the place where you type text

OpenStudy (anonymous):

Oh I know, but I mean I'm not sure how to draw out the question

OpenStudy (anonymous):

And that's all the information I have on it

OpenStudy (anonymous):

We just need to see the relationship the angles have to each other in order to solve the problem

OpenStudy (nali):

u mean it is a plain question with no picture involved above or under it that can help u in solving the problem

OpenStudy (anonymous):

just look at the diagram in the problem and give a basic representation

OpenStudy (anonymous):

you can't solve the problem without knowing how the angles are related

OpenStudy (anonymous):

Yeah, there is no picture. I don't know how to represent it or how to figure out if the angles are related or not.

OpenStudy (nali):

what does it say above the question?

OpenStudy (anonymous):

It's another question asking to name an alterior angle

OpenStudy (anonymous):

It has a picture. I'm not sure that they're related though.

OpenStudy (anonymous):

no they for sure are related just need to know how. try to draw the picture.

OpenStudy (anonymous):

It doesn't have to be good.

OpenStudy (anonymous):

I will attach it

OpenStudy (anonymous):

okay so angle 4 and 8 have the same value so set them equal to each other and solve for x

OpenStudy (anonymous):

3x=x+40 then you get 2x=40 so x=20

OpenStudy (anonymous):

so the angle must be 60

OpenStudy (anonymous):

Ohh. Okay. Thank you so much!

OpenStudy (anonymous):

no problem do you understand how I did that?

OpenStudy (anonymous):

No can you explain it?

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

okay can you tell that angle 4 and 8 must have the same value?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay so if they have the same value and you have the equations that represent that value for each of them then those equations have to be equal to each understand? (If not say so I will explain more)

OpenStudy (anonymous):

each other**

OpenStudy (anonymous):

A little

OpenStudy (anonymous):

alright so think about it like this now that you know each angle equals 60 imagine 60 next to each of the equations. So 60=3x and 60=x+40. Now since they both equal the same value then you can also say 3x=x+40. Follow so far?

OpenStudy (anonymous):

Ohh, yes, I do.

OpenStudy (anonymous):

alright =) so replace that 60 with another variable because we don't know the value at the beginning of the question. So angle = 3x angle =x+40 then we need to find x in order to find the answer and we do that by setting the equations equal to eachother and then solving for x. Once we solve for x we can plug it back into one of the equations to get our value for angle. know what I mean?

OpenStudy (anonymous):

Yes, I get it

OpenStudy (anonymous):

awesome =)

OpenStudy (anonymous):

Thank you so much! I appreciate all your help

OpenStudy (anonymous):

no problem dude =)